OpenStudy (anonymous):

Any answers: 10+X-(2-3X)=4 Solve for X

OpenStudy (saifoo.khan):

10 + x - 2 + 3x = 4

OpenStudy (anonymous):

x=-1

OpenStudy (saifoo.khan):

4x = -4

OpenStudy (saifoo.khan):

x = -1

OpenStudy (anonymous):

4x = -4 x = -1

OpenStudy (anonymous):

hi folks, if you do the problem in the parenthes?? first, then start at the beginnng and work your way to the right, -1 does not work. Or am I wrong?

OpenStudy (anonymous):

u mean when ur replacing value of x or when u need to find it

OpenStudy (anonymous):

Yes it does. 10+(-1)-(2-3(-1))=4 10-1-(2+3)=4 10-1-5=4 10-6=4 4=4

OpenStudy (anonymous):

Show how you did it so we can narrow down where the problem lies.

OpenStudy (anonymous):

I did the (2-3X) first, when replacing X with -1, this part of the problem becomes 2- -3 =-1. then 10 + -1=9- -1 = 8, not 4

OpenStudy (anonymous):

Inside the parenthesis you have (2-3x) Since x=-1 3x means 3*(-1) so 3x=-3 Now (2-3x) becomes (2-(-3)) So that is (2+3) So inside the parenthesis it's 5 The key seems to be: 2-(-3)=2+3=5 Two minuses give you a plus sign.

OpenStudy (radar):

Rudy's Right

OpenStudy (anonymous):

OK, I guess I am not understanding, Youhave a positive number 2 and you are taking away -3 from that gives you a negative number -1. Why?

OpenStudy (anonymous):

2-3=-1 But you don't have that. You have 2-(-3) and two negatives make a positive. So your: 2-(-3) is the same as 2+3 Which is 5 You had (2-3x) But the x=-1 So if you multiply 3*(-1) all it does is switch the sign of the 3. So from (2-3x) it goes to (2+3)

OpenStudy (radar):

2 - -3 -------- 5

OpenStudy (anonymous):

Thank you very much, I think I get it now. Thank you for your help.

OpenStudy (anonymous):

you're welcome.

OpenStudy (radar):

good job rudy, Glad you stayed with it, and now there is understanding.